OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points (-5/3, -1) and (-2, 9/2) ?

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

pooja195 (pooja195):

\[\LARGE \frac{9/2--1}{-2-5/2}=slope\] \[\LARGE\bf \frac{9/2+1}{-2-5/2}=slope\]

OpenStudy (anonymous):

5.5 / -4.5

OpenStudy (calculusxy):

\(\text{Slope} = \large \frac{y_2 - y_1}{x_2 - x_1}\) You have the coordinates (-5/3, -1) and (-2, 9/2) You can assign \(\frac{9}{2}\) to be \(y_2\) and \(-1\) to be \(y_1\) even though it's not a huge matter if you put either one. Then you can have \(-2\) to be \(x_2\) and \(-\frac{5}{3}\) to be \(x_1\). You will now have to substitute solve.

OpenStudy (anonymous):

-5.5 / -4.5 ??

OpenStudy (anonymous):

I divide them next right?

pooja195 (pooja195):

I'll let @calculusxy take over i don't want to confuse you :)

OpenStudy (calculusxy):

\[\large \frac{ \frac{ 9 }{ 2 } - (-1) }{ -2 - (-\frac{ 5 }{ 3 }) } = \frac{ \frac{ 11 }{ 2 } }{ -\frac{ 1 }{ 3 } }\]

OpenStudy (anonymous):

Im so confused...um what do I do next?

OpenStudy (calculusxy):

And then you have to divide. Remember that fractions are kind of like divisions the fraction bar indicates to divide in most circumstances. \[\large \frac{ 11 }{ 2 } \times -\frac{ 3 }{ 1 }\]

OpenStudy (calculusxy):

Solve for that and you should be able to get your slope.

OpenStudy (anonymous):

So I multiply 11/2 x -3/1????

OpenStudy (calculusxy):

yes

OpenStudy (calculusxy):

and what will you get?

OpenStudy (anonymous):

um, I got 16.5

OpenStudy (anonymous):

im soooo confused. sorry.

OpenStudy (anonymous):

I stopped at -5.5 / -4/5

OpenStudy (calculusxy):

So we would have: \[\large \frac{ 11 }{ 2 } \times -\frac{ 3 }{ 1 } = -\frac{ 33 }{ 2 } = -16.5\]

OpenStudy (calculusxy):

So your answer of 16.5 was partially correct, except that you had to put the negative sign in front!

OpenStudy (anonymous):

16.5 isnt a answer choice though...

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @calculusxy So your answer of 16.5 was partially correct, except that you had to put the negative sign in front! \(\color{#0cbb34}{\text{End of Quote}}\) \[\huge~\rm~-16.5\]

OpenStudy (anonymous):

These are the answer choices: 2/33, -2/33, -33/2, and 33/2

pooja195 (pooja195):

-16.5 in fraction form is -33/2

OpenStudy (anonymous):

Okay thanks.

OpenStudy (calculusxy):

\[\large \frac{ 11 }{ 2 } \times -\frac{ 3 }{ 1 } = -\frac{ 33 }{ 2 } = -16.5\] -33/2 is part of this (shown above)

OpenStudy (anonymous):

oh yeah, okay I see now.

OpenStudy (anonymous):

Thanks!

OpenStudy (calculusxy):

Welcome !